OpenStudy (anonymous):

Solve x + 3y = 9 3x – 3y = –13 *********************** (–1, 3 over 10) (1, negative 3 over 10) (–1, ten thirds) (1, negative ten thirds)

OpenStudy (anonymous):

is it -1 10/3?

OpenStudy (anonymous):

x+3y=9 solve for y in terms of x 3y=9-x y=3-(1/3)x plug this into second equation and solve for x 3x-3(3-(1/3)x))=-13 3x-9+x=-13 4x-9=-13 4x=-4 x=-1 plug back into first equation and solve for y -1+3y=9 3y=10 y=10/3